OpenStudy (anonymous):

Algebra Help, Modeling with quadratic functions, writing an equation of a parabola???

OpenStudy (anonymous):

3 points, (3,-6), (1,-2), and (6,3)

OpenStudy (anonymous):

\[y=ax^2+bx+c\]

OpenStudy (anonymous):

\[-6=a(3)^2+b(3)+c\] \[-2=a(1)^2+b(1)+c\] \[3=a(6)^2+b(6)+c\]

OpenStudy (anonymous):

\[9a+3b+c=-6\] \[a+b+c=-2\] \[36a+6b+c=3\]

OpenStudy (anonymous):

I don't get what to do after this part :(